### v1

# Rich Semantic Knowledge Enhanced Large Language Models for Few-shot Chinese Spell Checking

Abstract
Chinese Spell Checking (CSC) is a widely used technology, which plays a vital role in speech to text (STT) and optical character recognition (OCR). Most of the existing CSC approaches relying on BERT architecture achieve excellent performance. However, limited by the scale of the foundation model, BERT-based method does not work well in few-shot scenarios, showing certain limitations in practical applications. In this paper, we explore using an in-context learning method named RS-LLM (Rich Semantic based LLMs) to introduce large language models (LLMs) as the foundation model. Besides, we study the impact of introducing various Chinese rich semantic information in our framework. We found that by introducing a small number of specific Chinese rich semantic structures, LLMs achieve better performance than the BERT-based model on few-shot CSC task. Furthermore, we conduct experiments on multiple datasets, and the experimental results verified the superiority of our proposed framework.
Rich Semantic Knowledge Enhanced Large Language Models for Few-shot Chinese Spell Checking
[…]
1 Introduction
Spell checking (SC) aims to utilize intelligent methods to automatically identify and correct errors in text. This technology facilitates nature language processing applications to correct the errors from different text input systems, such as speech to text (STT) and optical character recognition (OCR). In recent years, SC has attracted tremendous attention from the research community Chodorow et al. (2007); Malmi et al. (2019); Mallinson et al. (2020). Chinese spell checking (CSC) specifically refers to SC for Chinese text. Compared with the relatively complete SC technology, CSC still cannot be perfectly applied to various practical scenarios, and there are still many problems that need to be solved Zhang et al. (2023).
As an ideogram, the usage and structure of Chinese are very different from English, which leads to different challenges for CSC and SC. First of all, the pronunciation of Chinese varies greatly, and it is difficult to easily infer the glyphs from the pronunciation. When listening to a piece of Chinese speech, if you do not understand the context, the text result obtained based on the speech is likely to contain a large number of homophones. In addition, the glyph structure of Chinese is more diverse, resulting in more types of errors. Therefore, CSC task mainly needs to handle two types of error texts. One is a text based on the STT system, which contains a large number of homophonic errors. The other is text generated by OCR-based systems that mainly contains glyph errors. To address these two types of spelling errors, most of the existing studies use models based on BERT architecture, then introducing the external glyph-phonetic features Ji et al. (2021); Xu et al. (2021); Ji et al. (2021); Li et al. (2022).
In practical Internet applications, the catchwords from hot topics vary rapidly and unlabeled incorrect sentences emerge constantly, resulting in few-shot scene for CSC. However, existing BERT-based models are difficult to be conduct in few-shot scene because of the limited scale of the foundation model. Large language models (LLMs) show remarkable ability on semantic analyzing, positioning them to become an optimal foundation model for CSC. This paper focus on CSC in few-shot scene. We build a Chinese rich semantic corpus (See details in Section 4.1). Besides, we choose LLMs as foundation and integrate Chinese rich semantic knowledge by in-context learning. Furthermore, we conduct experiments on several datasets. The contributions of this study are summarized as follows:
  • We propose an in-context learning based method to introduce LLMs to CSC task, which improves the performance of few-shot scenarios.
  • We propose the paradigm of prompt template designing for CSC.
  • We conduct experiment to compare different Chinese Rich Semantic structures. And we find the best structures for LLMs based CSC tasks.
Figure 1: The performce of CSC tasks using LLM (ChatGPT) without specific restrictions on input prompts.
2 Related Work
Due to the lack of parallel corpus training data, early CSC methods mainly rely on linguistic knowledge to manually design rule-based methods Mangu and Brill (1997); Jiang et al. (2012). Subsequently, machine learning models become the main paradigm for CSC tasks Chen et al. (2013); Yu and Li (2014). Machine learning typically employs language models, such as n-grams, to detect error locations. Then uses confusion sets and character similarities to correct potential misspelled characters and candidate correct characters, and finally scores replace sentences through the language model to determine the best correction solution Liu et al. (2013); Xie et al. (2015).
The field of CSC advances significantly with the development of deep learning, particularly through pre-trained models like BERT Devlin et al. (2019). Pre-trained models such as BERT are known for their context awareness and transfer learning capabilities. Most current CSC models with better performance use BERT as the baseline model. Hong et al. (2019) innovates by modeling CSC as a BERT token classification task, utilizing the Confidence-Character Similarity Decoder (CSD). Zhang et al. (2020) enhances this approach by combining error identification and correction losses with a soft mask strategy. Addressing the prevalent issue of phonetic and glyph similarities in spelling errors, the integration of these features with semantic information is now a primary research focus. Liu et al. (2021) suggests incorporating confusion sets into pre-training with a GRU network to better mimic real errors and model character sound-shape similarities. Xu et al. (2021) proposes a multi-modal approach to capture semantic, phonetic, and graphical information and the use of adaptive gating modules to merge semantic, phonetic, and glyph features in CSC. Ji et al. (2021) introduces SpellBERT, which integrates radical features into character representation using a graph convolution network. The SCOPE model Li et al. (2022) further develops this field by adding pronunciation prediction tasks in training, forging a deeper connection between semantics and phonetics, and employing iterative reasoning strategies to bolster CSC model performance.
Recently, the development and advancement of LLMs have brought natural language processing to the next stage. Li et al. (2023b) analyzes the correction ability of the OpenAI [^1] https://openai.com/’s existing LLMs and finds that they still fall short of the CSC capabilities of previous fine-tuned models.
3 Preliminary
3.1 Chinese Rich Semantics
Chinese is the carrier of the inheritance and development of Chinese civilization. As a hieroglyphic script that has lasted for thousands of years, Chinese characters have rich semantic structures, including traditional characters, glyphs, phonetic, tones and other features. Chinese contains a large number of phonograms, whose pronunciation and meaning are contained in radicals. For example, "海 (sea)", "河 (river)" and "湖 (lake)" all have the radical "氵", related to water Sun et al. (2021). For another example, "株 (zhū)", "诛 (zhū)" and "珠 (zhū)" all have the same phonetic tone "朱 (zhū)", so the pronunciation of these characters is similar. In essence, radicals can be regarded as a kind of classification label information. This classification information serves as traditional context. The additional supplement of semantic features has important semantic value. In addition to the features at the glyph level, Chinese also has features such as morpheme diversity, ambiguity, and structural diversity.
3.2 Task Definition
Formulate definition of CSC: CSC focuses on identifying and rectifying spelling errors in a given text sequence. Essentially, it involves processing a sequence of characters $X=\{x_{1},x_{2},...,x_{n}\}$ and producing a corrected sequence $Y=\{y_{1},y_{2},...,y_{n}\}$, where $n$ is the number of words. Here, $X$ denotes the initial text with potential errors, and $Y$ signifies the amended correct text. The two sequences $X$ and $Y$ have the same length.
Few-shot Learning on CSC: Few-shot learning refers to provide a model named $L$ with $K$ pairs of contexts and corresponding required answers as examples, and then a context that requires model $L$ to reason. The goal of model $L$ is to generate an appropriate answer for the last context. During inference, model $L$ is guided by $K$ context-answer pairs without any updates to its parameters Brown et al. (2020).
In few-shot learning for CSC task, we set $K$ to 3. Then we have a few-shot learning prompt $P$, which contains three sentences pairs:
$P=\{p_{1},p_{2},p_{3}\}$ (1)
where $p_{i}$ contains an incorrect sentence and its corresponding correct sentence.
Then the sentence $X$ that need to be spell checked:
$X=\{x_{1},x_{2},\ldots,x_{n}\}$ (2)
where $x_{i}$ denotes a word of sentence $X$. $n$ is the length of sentence $X$.
For model $L$, the inputs are $P$ and $X$, and the output is the correct sentence $Y$ corresponding to the sentence $X$.
$Y=L(P,X)$ (3)
$Y=\{y_{1},y_{2},\ldots,y_{n}\}$ (4)
where $y_{i}$ denotes a word of sentence $Y$ and $n$ is the length of sentence $Y$.
3.3 Challenges of LLMs based CSC Tasks
The versatility of LLMs gives them significant text polishing capabilities. Since there are no specific restrictions on input prompts, LLMs tend to perform freely in CSC tasks. However, free play of LLMs may result in LLMs outputting sentences that are completely grammatically correct. However, the sentences output by LLMs are different from the standards established by existing CSC datasets and evaluation indicators. Therefore, these existing traditional datasets are used without specific restrictions on input prompts. It becomes challenging to objectively and realistically evaluate the spell checking performance of LLMs. An example of using LLM (ChatGPT) to perform a CSC task without setting specific restrictions on input prompts is shown in Fig.1. From Fig.1 it can be observed that two main problems are prone to occur in LLM when performing CSC tasks, one is the length of the input sentence and the output sentence are inconsistent. For instance, the sentence "I am really bald (我真胡秃)", which should be corrected to "I am really confused (我真糊涂)". However, LLM correct this sentence to "I am really sorry (我真的抱歉)". Another problem is that LLM easily rewrites the input sentences semantically. For example, "I hope you have a good time dancing without (希望你们好好的跳无)", in which "dancing without (跳无)" should be corrected to "dancing (跳舞)". Instead, LLM correct this sentence to "I hope you have a good New Year (希望你们好好的过年)".
Figure 2: Task-specific few-shot prompt for CSC tasks. We marked the semantic structure information (speech and radicals) and key information related to the task features in the prompts in different color fonts.
4 Method
4.1 Chinese Rich Semantic Corpus
Chinese, a logo-graphic language, inherently possesses a rich semantic depth in its character glyph, potentially enhancing the expressiveness of LLMs. Our research focuses on the GB2312 simplified Chinese coding table, a standard set by China’s State Administration of Standards on May 1, 1981222https://openstd.samr.gov.cn/bzgk/gb. This table consists of 6,763 Chinese characte rs, divided into 3,755 primary and 3,008 secondary characters. It contains most commonly used Chinese characters. In order to obtain the detailed information of this coding table, we collected various basic attributes of each Chinese word, such as its phonetic, radical, phonetic notation, strokes, outside strokes (the strokes except the radical), stroke order, structure, Unicode, Wubi code, Cangjie code, Zheng code, Four-corner code, as well as glyph images from different historical periods. Despite the large amount of data collected, we note issues with the quality and completeness of the data. To address these issues, we manually annotate the collected information to ensure a more accurate and comprehensive dataset. We give the attributes of the word 海 (sea) in the dataset as shown in Fig.3. In order to understand and better utilize these properties, we classify these features into the following three categories:
  • Phonetic Features:Phonetic which uses Latin letters to represent the pronunciation of Chinese characters. Phonetic notation is a phonetic system that uses symbols to represent the sounds of Chinese characters. Zheng Code is a Chinese character input method that assigns codes based on the initials of the phonetic pronunciation.
  • Glyph Features:Radical is a category according to the type and side of the Chinese characters, and all the Chinese characters are bound to be classified in a certain radical. Strokes refers to the number of lines needed to write Chinese character. Outside strokes mean the number of lines needed to write Chinese character except the radical. Structure refers to the internal organization of Chinese characters, including the arrangement of radicals, strokes, and components.
  • Input Coding Features:Stroke Order indicates the sequence in which strokes are written when forming a Chinese character. Proper stroke order is important for correct character writing. Cangjie code assigns codes to characters based on their shapes and components. Unicode is an international standard for character encoding that assigns unique codes to characters from various writing systems, including Chinese characters. Wubi code assigns codes based on the five basic strokes. Four-Corner code is a Chinese character input method that assigns codes based on the shapes of the four corners of a character.
Figure 3: Various attributes of the word 海 (sea) in the Chinese Rich Semantics.
4.2 In Context Learning for CSC
A variety of studies Xie et al. (2022); Dai et al. (2023); Bansal et al. (2023) have revealed that LLMs exhibit exceptional in-context learning capabilities Dong et al. (2023). In-context learning can quickly improve the task-specific performance of LLMs by providing them with a limited set of task-related examples, which requires extremely high computational resources and is extremely expensive and difficult to fine-tune LLMs. In-context learning can quickly adapt to most LLMs without the need for separate training for each LLM. As mentioned in Section 3.3, when LLMs perform CSC tasks without specific task constraints on input prompts, the answers given by LLMs are very likely to be inconsistent with the existing evaluation indicators of CSC tasks. Therefore, in order to accurately and objectively explore the performance ability of LLMs on CSC tasks, we design the task-specific prompts as shown in Fig.2.
First of all, we give the identity and task description of LLMs in the prompt, according to Li et al. (2023a), we know that giving LLMs emotional motivation (you are an excellent Chinese Spell Checking model) can effectively stimulate LLM’s comprehension. The task description requires the LLMs to only correct spelling mistakes, thus limiting the LLM’s semantic rephrasing of the input sentence. Since CSC task requires the input sentence and the output sentence to be equal in length, the LLMs is required not to add any other explanations and descriptions when performing CSC task, so as to require that the length of the sentence output by the LLMs and the input error correction sentence are equal. Please note that this does not mean that LLMs will not semantically rewrite the input sentence or change the sentence length after such a request. This can only minimize the role of LLMs in these issues.
Next, We give three pairs of input sentences and their corrected sentences, carefully selected from the corresponding training set of the dataset we use. From Liu et al. (2010), Chinese text errors are primarily caused by characters that are visually and phonetically similar. These three sentence pairs contain a set of phonetic errors, a set of glyph errors, and a set of correct sentences that do not need to be corrected.
We add specific task requirements to the input sentences of three pairs of sentences, which is different from the previous task description. We concretize the task here, asking LLMs to find spelling errors in sentences and replace wrong words with correct words. This allows LLMs to have a clearer understanding of CSC tasks. Next, we append the length of the input sentence to the end to indicate the LLMs that we limit the length of the output sentence in the first prompt. We also include the phonetic and radical information of each Chinese character in the input sentence, drawn from Chinese Rich Semantic Corpus outlined in Section 4.1. It is hoped that by adding this information to the prompt, the LLMs can strengthen their understanding of the input sentence. Therefore, LLMs can better correct phonetic and visual errors that may occur in the sentences. These are in the form of four pairs of historical dialogues that form a few-shot prompt input to LLMs, hoping to stimulate the in-context learning capabilities of LLMs at once.
For the sentences to be corrected in the test set, we use the prompt of the input sentence in the few-shot prompt to construct the same prompt for the sentences to be corrected in the testset and then enter LLMs to hope that the LLMs will correct the spelling of the error corrected sentences through in-context learning. Considering the length limit of LLMs on the length of the input (including historical dialogue), and the deterioration of the semantic understanding of LLMs with long input lengths. Before spell-checking each sentence, we clear the historical memory of the LLMs and add our few-shot prompt. In order to verify that the prompt we proposed is effective, we counted the number of sentences whose length remained unchanged before and after using a specific prompt and not using a specific prompt. The results are in Table 1. Experiments were performed using the API of GPT-3.5-Turbo.
Data without prompt
SIGHAN15 578 724
LAW 83 125
MED 237 262
ODW 294 347
Table 1: The number of sentences whose length does not change using specific prompts and not using specific prompts on different data sets
5 Experiments
5.1 Datasets and Evaluation Metrics
Datasets: In the evaluation of CSC task, we use two highly used datasets. The first, SIGHAN15 Tseng et al. (2015), consists of handwritten samples from learners of Chinese as a Second Language (CSL). These samples provide a rich source of real-world language usage by non-native speakers, offering insights into common errors and patterns in learning Chinese. The second dataset Lv et al. (2023) is specialized and segmented into three distinct domains to provide a more comprehensive understanding of language use in specific contexts. For the LAW domain, data is sourced from the stems and choices of multiple-choice questions in judicial examinations, reflecting the formal and technical language of the legal field. The MED domain encompasses data from question-and-answer pairs drawn from online medical consultations, showcasing the specific terminology and communication style in healthcare. The ODW (Official Document Writing) domain includes data from various official documents such as news, policies, and state reports on national conditions, representing formal and structured writing styles. The statistics of the test data from these four datasets we used are shown in Table 2.
Test Data #Sent Avg. Length #Errors
SIGHAN15 1100 30.6 703
LAW 500 29.7 356
MED 500 49.6 345
ODW 500 40.5 403
Table 2: Statistical information regarding the dataset in our experiments, which includes the total count of sentences (#Sent), the average length of these sentences (Avg. Length), and the total number of spelling mistakes (#Errors).
Evaluation Metrics: Following the previous work Xu et al. (2021); Lv et al. (2023), we evaluate the performance of CSC at two levels: detection and correction. At the detection level, a sentence is considered correct if it successfully identifies all spelling errors. At the correction level, the model not only identifies but also rectifies all erroneous characters by replacing them with the correct ones. We provide accuracy, precision, recall, and F1 scores for both evaluation levels.
5.2 Baselines
In order to explore whether Chinese Rich Semantics in LLMs we proposed can be improved on CSC task, we choose two widely foundational models.
gpt-3.5-turbo [^3] https://platform.openai.com/docs/api: One of a series of LLMs provided by OpenAI, which can be accessed through the API. The gpt-3.5-turbo serves as the underlying module for ChatGPT and is trained on GPT-3 using Reinforcement Learning from Human Feedback (RLHF).
glm-3-turbo [^4] https://open.bigmodel.cn/dev/api#glm-3-turbo: An open-source bilingual (Chinese-English) chat model. The glm-3-turbo is a variant of ChatGLM3 that is specifically optimized for dialogue generation tasks, and it has also achieved excellent results in various natural language processing tasks.
To evaluate the performance of LLMs, we choose advanced CSC models as our baseline references:
BERT Devlin et al. (2019): The basic BERT encodes the input sentence to get semantic information, followed by using a classifier to select the correct character from the vocabulary.
ChineseBERT Sun et al. (2021): The basic ChineseBERT encodes the input sentence to get semantic, phonetic, and graphical information, then use a classifier to select the correct character from the vocabulary.
ReaLiSe Xu et al. (2021): The CSC model captures semantic, phonetic, and graphical information of input characters using multimodality for prediction.
5.3 Implementation Details
In all CSC models in comparative experiments, the parameters are set to the optimal parameters in the open source code corresponding to the model. In order to achieve few-shot scenarios, all comparative experiments randomly selected 10 pieces of data from the training set, trained for 1000 rounds, and then selected the round with the best performance. The LLMs experiment uses the same few-shot prompt designed for CSC task as RS-LLMs, but does not add semantic information. In gpt-3.5-turbo-related experiments, we set the temperature to 0.8. In glm-3-turbo-related experiments, we set the temperature to the default 0.95. All experiments were conducted on a RTX-4090 with 24G memory.
Dataset Method Detection Level Correction Level
Acc. Pre. Rec. F1 Acc. Pre. Rec. F1
SIGHAN15 BERT 17.8 15.1 16.1 15.8 15.4 3.6 2.4 2.8
ChineseBERT 16.2 10.4 12.9 11.5 13.2 2.7 5.5 3.6
ReaLiSe 27.5 16.2 18.6 17.3 22.3 3.6 3.1 3.4
glm-3-turbo 45.6 28.9 40.1 33.6 42.2 23.9 33.1 27.7
RS-glm-3-turbo 46.4 29.7 40.1 34.2 43.5 25.3 34.2 29.1
gpt-3.5-turbo 37.0 24.1 38.6 29.7 35 21.6 34.6 26.6
RS-gpt-3.5-turbo 46.0 27.3 36.0 31.1 44.6 25.1 33.1 28.5
MED BERT 14.6 7.1 13.3 9.2 12.6 1.8 0.4 0.6
ChineseBERT 14.0 8.1 10.4 9.1 11.2 0.6 1.5 0.9
ReaLiSe 26.2 9.7 16.8 12.3 18.6 1.5 0.9 1.1
glm-3-turbo 57.0 34.7 49.6 40.8 52.6 27.9 39.8 32.8
RS-glm-3-turbo 59.4 37.4 47.4 41.8 56.0 31.5 39.8 35.2
gpt-3.5-turbo 30.8 17.5 32.3 22.7 28.8 15.1 27.9 19.6
RS-gpt-3.5-turbo 39.0 20.7 34.1 25.8 36.4 17.2 28.3 21.4
LAW BERT 14.4 8.8 12.3 10.3 12.7 1.35 0.4 0.6
ChineseBERT 15.0 12.3 14.3 13.2 13.5 0.8 1.6 1.1
ReaLiSe 23.3 15.4 18.2 16.7 20.2 1.6 1.5 1.5
gpt-3.5-turbo 46.6 31.9 45.9 37.6 43.8 28.1 40.4 33.1
RS-gpt-3.5-turbo 56.2 37.6 43.9 40.5 54.0 33.9 39.6 36.5
ODW BERT 16.8 13.2 15.6 14.3 13.1 4.4 1.5 2.3
ChineseBERT 15.9 12.2 14.2 13.1 11.6 2.0 3.8 2.7
ReaLiSe 30.2 18.5 26.8 21.9 25.4 5.4 4.2 4.7
gpt-3.5-turbo 65.0 52.3 66.0 58.0 61.8 47.4 59.9 52.9
RS-gpt-3.5-turbo 75.6 63.7 62.2 62.9 73.4 58.4 58.0 58.7
Table 3: The performance of all baselines and RS-LLMs. RS-gpt-3.5-turbo means RS-LLM on gpt-3.5-turbo and RS-glm-3-turbo means RS-LLM on glm-3-turbo. The glm-3-turbo and gpt-3.5-turbo utilizing identical few-shot prompts as RS-LLM, but without semantic information.
Figure 4: Detect-F1 score trends on the test dataset.
Figure 5: Correct-F1 score trends on the test dataset.
5.4 Experimental Results
Table 3 shows the evaluation results of our RS-LLM against other models on the SIGHAN15, LAW, MED and ODW test sets. We found that the experimental results of LLM have certain fluctuations with temperature changes, so our baseline experiments of LLM and RS-LLM experiments on the same model all use the same temperature. It is observed that our RS-LLM outperforms all baselines on all metrics in most cases for all datasets. In particular, LLM uses the same few-sample cues as RS-LLM but does not add semantic information when performing the CSC task. The experimental results of this paper verify the effectiveness and superiority of LLM semantic information in CSC tasks.
As shown in the table reftable:result, our spell checking method, RS-LLM over glm-3-turbo, achieves 0.6% error detection and 2.1% correction on SIGHAN15 dataset using Rich Semantic compared to standard LLM. And RS-LLM on glm-3-turbo registers 1% error detection and 2.4% correction rate were achieved on the MED dataset. RS-LLM on gpt-3.5-turbo, achieved a 1.4% correction rate of 1.4% error detection on SIGHAN 15 using Rich Semantic. In the LAW dataset, RS-LLM on gpt-3.5-turbo beats gpt-3.5-turbo by 2.8% in detection and 2.6% in correction, while RS-LLM on gpt-3.5-turbo outperforms gpt-3.5-turbo by 3.1% and 1.8% in the MED dataset. On the ODW dataset, our method also shows significant gains, with RS-LLM on gpt-3.5-turbo leading gpt-3.5-turbo by 4.9% in detection and 5.8% in correction.
In addition, it’s clear that the CSC model based on BERT performs worse than LLM in few-shot scenarios, both at the detection level and the correction level. The ability of the CSC model based on BERT to detect and correct incorrect characters excessively relies on the training data of CSC, especially the ability to correct incorrect characters.
5.5 Analyses and Discussions
5.5.1 Impact of Different In-context Learning Approaches
Fig.4 and Fig.5 show the evaluation results of our RS-LLM and LLM with different context learning methods. For specific experimental results, see the table 4 in the appendix 5. From Fig.4 and Fig.5, we find that the performance of the CSC task on LLM improves on most datasets as the number of examples in the context increases, both in terms of detection and correction. Furthermore, RS-LLM is more effective than LLM in few-shot scenarios compared to zero-shot and one-shot scenarios. Such experimental results reflect the effectiveness of our in-context learning approach designed for CSC task.
5.5.2 Impact of Different Semantic Information
In order to further study the impact of semantic information on LLMs when performing CSC tasks. We design to use phonetic information, radical information, structural information and strokes information. Each of these four types of information is added to our zero-shot prompt, one-shot prompt, and few-shot prompt. To explore the impact of these four prompts on the test set of SIGHAN15, LAW, MED, and ODW in three different contexts of learning. The specific experimental results is Table 5 in Appendix 5.
Experimental results show that single speech, radical, structure and stroke information can improve CSC task on most datasets in few-shot scenarios. It is worth noting that speech and radical information are basically enhanced on all datasets, and structural and stroke information can improve the performance of LLM on the CSC task on some datasets. But this improvement is relatively small compared to other information.
6 Conclusion and Future Work
In this paper, we introduce an in-context learning method named RS-LLM for CSC task, one of whose core components is using Chinese Rich Semantics in LLMs for CSC task. RS-LLM utilizes adding a small number of specific Chinese rich semantic information into a specific few-shot prompt set for CSC task, aiming to allow LLMs to have a fuller understanding of the semantics when doing CSC task. Experimental results show that this LLMs-independent approach can help existing LLMs better recognize and correct phonetically and visually erroneous characters in CSC tasks. Considering the impact of the similarity of errors in the few-shot prompt and errors in the sentence on LLMs’ understanding of the sentence when LLMs perform CSC task. Next we will try to enhance our in-context learning method through similarity retrieval when constructing few-shot prompts.
Limitations
To verify the effectiveness of RS-LLM, we conducted extensive experiments on two benchmark datasets of different domains and scales. The results indicate that RS-LLM delivers SOTA results in few -shot scenarios. However, since most of the errors in the CSC dataset are attributed to visual and phonetical errors, we incorporate phonetic and radical information into the prompt template. However, it is difficult to ensure that the existing manually formulated prompt templates are optimal, and the optimal prompt sentences for CSC require further research. Furthermore, relevant examples have been carefully selected to enable LLMs to identify potential visual and speech errors in a small number of scenarios. No consideration was given to the ability to motivate the LLMs through semantic similarity. We recognize these two limitations and plan to address them in future research efforts.
Ethics Statement
We adhere to and advocate for the principles outlined in the ACL Code of Ethics. The primary focus of this paper is on the task of CSC, with an objective to enhance the performance of LLMs in this task by incorporating semantic knowledge into the template. The datasets utilized in our research are obtained from openly published sources, ensuring they are free from privacy or ethical concerns. In our approach, we consciously avoid introducing or magnifying any social or ethical biases in the model or data. Consequently, we anticipate no direct social or ethical challenges as a result of our research.
Acknowledgments
Appendix A The Detailed Experimental Results
In-context Learning Approaches Dataset Method Detection Level Correction Level
Acc. Pre. Rec. F1 Acc. Pre. Rec. F1
zero-shot SIGHAN15 glm-3-turbo 47.1 28.7 29.4 29.0 44.1 22.7 23.3 23.0
RS-glm-3-turbo 47.9 26.5 27.7 27.1 45.8 22.4 23.5 22.9
gpt-3.5-turbo 33.6 22.7 36.6 28.1 31.8 20.4 32.9 25.2
RS-gpt-3.5-turbo 42.5 25.8 34.9 29.7 40.4 22.8 30.9 26.2
LAW glm-3-turbo 36.4 20.7 26.7 23.3 34.4 17.6 22.8 19.9
RS-glm-3-turbo 60.8 42.7 40.4 41.5 57.6 36.1 34.1 35.1
gpt-3.5-turbo 28.4 20.6 34.1 25.7 26.4 18.2 30.2 22.7
RS-gpt-3.5-turbo 44.8 26.4 32.2 29.0 42.6 22.8 27.8 25.1
MED glm-3-turbo 49.0 26.9 35.4 30.6 46.6 22.9 30.1 26.0
RS-glm-3-turbo 53.4 26.5 30.9 28.6 51.2 22.4 26.1 24.1
gpt-3.5-turbo 33.8 19.6 34.9 25.1 31.6 16.9 30.1 21.6
RS-gpt-3.5-turbo 36.6 17.8 28.3 21.9 33.6 13.7 21.7 16.8
ODW glm-3-turbo 54.8 45.8 30.9 36.9 52.4 39.0 26.3 31.4
RS-glm-3-turbo 69.6 58.1 45.0 50.8 66.6 50.7 39.3 44.3
gpt-3.5-turbo 38.8 31.3 50.7 38.7 35.6 27.5 44.7 34.1
RS-gpt-3.5-turbo 57.2 41.2 43.5 42.3 53.6 34.7 36.6 35.6
one-shot SIGHAN15 glm-3-turbo 43.7 28.4 40.1 33.3 39.8 22.8 32.2 26.7
RS-glm-3-turbo 42.8 23.6 29.8 26.3 40.6 20.0 25.1 22.2
gpt-3.5-turbo 29.6 21.2 36.6 26.9 27.7 18.9 32.7 24.0
RS-gpt-3.5-turbo 38.6 23.2 33.1 27.3 36.6 20.4 29.2 24.1
LAW glm-3-turbo 61.6 47.7 60.0 53.1 57.2 40.8 51.4 45.5
RS-glm-3-turbo 60.8 43.1 48.6 45.7 57.8 37.9 42.8 40.2
gpt-3.5-turbo 34.6 24.8 40.4 30.8 32.2 21.9 35.7 27.2
RS-gpt-3.5-turbo 41.6 24.3 32.2 27.7 39.6 21.3 28.2 24.3
MED glm-3-turbo 52.6 31.8 46.9 37.9 50.0 27.9 41.2 33.3
RS-glm-3-turbo 52.8 29.7 39.4 33.8 49.2 23.7 31.4 27.0
gpt-3.5-turbo 16.0 11.8 24.8 16.0 14.2 9.9 20.8 13.5
RS-agpt-3.5-turbo 25.4 13.8 25.2 17.8 24.0 12.1 22.1 15.7
ODW glm-3-turbo 80.0 69.5 68.7 69.1 76.4 62.6 61.8 62.2
RS-glm-3-turbo 74.2 62.5 54.6 58.3 70.2 53.7 47.0 50.1
gpt-3.5-turbo 40.8 31.9 50.0 38.9 38.6 29.2 45.8 35.7
RS-gpt-3.5-turbo 68.2 53.4 48.1 50.6 65.6 47.9 43.1 45.4
few-shot SIGHAN15 glm-3-turbo 45.6 28.9 40.1 33.6 42.2 23.9 33.1 27.7
RS-glm-3-turbo 46.4 29.7 40.1 34.2 43.5 25.3 34.2 29.1
gpt-3.5-turbo 37.0 24.1 38.6 29.7 35 21.6 34.6 26.6
RS-gpt-3.5-turbo 46.0 27.3 36.0 31.1 44.6 25.1 33.1 28.5
LAW glm-3-turbo 64.4 56.8 49.0 52.6 62.0 51.4 44.3 47.6
RS-glm-3-turbo 64.6 56.3 49.0 52.4 61.8 50.5 43.9 47.0
gpt-3.5-turbo 46.6 31.9 45.9 37.6 43.8 28.1 40.4 33.1
RS-gpt-3.5-turbo 56.2 37.6 43.9 40.5 54.0 33.9 39.6 36.5
MED glm-3-turbo 57.0 34.7 49.6 40.8 52.6 27.9 39.8 32.8
RS-glm-3-turbo 59.4 37.4 47.4 41.8 56.0 31.5 39.8 35.2
gpt-3.5-turbo 30.8 17.5 32.3 22.7 28.8 15.1 27.9 19.6
RS-gpt-3.5-turbo 39.0 20.7 34.1 25.8 36.4 17.2 28.3 21.4
ODW glm-3-turbo 81.0 69.9 69.1 69.5 76.6 61.4 60.7 61.0
RS-glm-3-turbo 78.0 67.2 64.1 65.6 73.4 58.0 55.3 56.6
gpt-3.5-turbo 65.0 52.3 66.0 58.0 61.8 47.4 59.9 52.9
RS-gpt-3.5-turbo 75.6 63.7 62.2 62.9 73.4 58.4 58.0 58.7
Table 4: The performance of different In-context Learning approaches added in prompts on LLMs. RS-gpt-3.5-turbo means RS-LLM on gpt-3.5-turbo and RS- glm-3-turbo means RS-LLM on glm-3-turbo.
Dataset Method Detection Level Correction Level
Acc. Pre. Rec. F1 Acc. Pre. Rec. F1
SIGHAN15 gpt-3.5-turbo 37.0 24.1 38.6 29.7 35.0 21.6 34.6 26.6
gpt-3.5-turbo+拼音 43.8 26.3 35.5 30.2 42.2 23.8 32.2 27.4
gpt-3.5-turbo+部首 44.3 25.7 34.2 29.3 41.5 22.9 30.5 26.2
gpt-3.5-turbo+结构 40.1 20.7 26.1 23.1 38.6 18.2 22.9 20.3
gpt-3.5-turbo+笔画 43.0 24.7 32.0 27.9 40.6 20.8 27.0 23.5
RS-gpt-3.5-turbo 46.4 30.0 42.0 35.0 43.7 26.1 36.6 30.5
LAW gpt-3.5-turbo 46.6 31.9 45.9 37.6 43.8 28.1 40.4 33.1
gpt-3.5-turbo+拼音 54.6 35.3 42.8 38.7 52.6 32.0 38.8 35.1
gpt-3.5-turbo+部首 53.8 34.9 42.8 38.5 52.0 32.1 39.2 35.3
gpt-3.5-turbo+结构 54.0 35.8 43.9 39.4 52.0 32.6 40.0 35.9
gpt-3.5-turbo+笔画 51.4 34.3 45.5 39.1 48.8 30.5 40.4 34.7
RS-gpt-3.5-turbo 56.2 37.6 43.9 40.5 54.0 33.9 39.6 36.5
MED gpt-3.5-turbo 30.8 17.5 32.3 22.7 28.8 15.1 27.9 19.6
gpt-3.5-turbo+拼音 35.4 18.5 31.4 23.3 33.4 15.9 27.0 20.0
gpt-3.5-turbo+部首 34.0 17.7 30.5 22.4 32.0 15.1 26.1 19.2
gpt-3.5-turbo+结构 36.6 19.6 33.2 24.7 34.4 16.8 28.3 21.1
gpt-3.5-turbo+笔画 23.0 12.8 24.3 16.7 21.6 11.1 21.2 14.6
RS-gpt-3.5-turbo 39.0 20.7 34.1 25.8 36.4 17.2 28.3 21.4
ODW gpt-3.5-turbo 65.0 52.3 66.0 58.0 61.8 47.4 59.9 52.9
gpt-3.5-turbo+拼音 71.8 57.5 57.3 57.3 69.2 52.5 52.3 52.4
gpt-3.5-turbo+部首 73.0 59.5 58.4 58.9 70.4 54.5 53.4 53.9
gpt-3.5-turbo+结构 71.0 55.6 59.2 59.2 67.8 49.8 53.1 51.4
gpt-3.5-turbo+笔画 73.4 59.2 60.3 59.7 70.4 53.6 54.6 54.1
RS-gpt-3.5-turbo 75.6 63.7 62.2 62.9 73.4 58.4 58.0 58.7
Table 5: The performance of different semantic information is added to few-shot prompts on LLMs. ’拼音’ means phonetic information, ’部首’ means radical information, ’结构’ means structural information, ’笔画’ means strokes information.
In this section, we introduce the specific experimental results of different in-context learning methods under different semantic information.